Find three consecutive integers whose squares add to 110.
Answer by checkley71(8403)

X^2+(X+1)^2+(X+2)^2=110 OR X^2+X^2+2X+1+X^2+4X+4=110 OR 3X^2+6X+5=110 OR
3X^2+6X-105 OR X^2+2X-35=0 OR (X+7)(X-5)=0 OR X+7=0 OR X=-7 & X-5=0 OR X=5 THEN
X+1=6 & X+2=7 THUS 5, 6 & 7 ARE THE THREE CONSECUTIVE INTERGERS WHEN SQUARED EQUAL 110
PROOF 5^2+662+7^2=110 OR 25+36=49=110 OR 110=110